### v1

# Enhanced SPS Velocity-Adaptive Scheme: Access Fairness in 5G NR V2I Networks

Abstract
Vehicle-to-Infrastructure (V2I) technology enables information exchange between vehicles and road infrastructure. Specifically, when a vehicle approaches a roadside unit (RSU), it can exchange information with the RSU to obtain accurate data that assists in driving. As the 3rd Generation Partnership Project (3GPP) Release 16, which includes the 5G New Radio (NR) Vehicle-to-Everything (V2X) standards, vehicles typically adopt mode-2 communication using sensing-based semi-persistent scheduling (SPS). By this approach, vehicles identify resources through a selection window and exclude ineligible resources based on information from a sensing window. However, vehicles often drive at different speeds, resulting in varying amounts of data transmission with RSUs as they pass by, which leads to unfair access. Therefore, developing an scheme that accounts for different vehicle speeds to achieve fair access across the network is essential. This paper formulates an optimization problem for vehicular environment and proposes a multi-objective optimization scheme to address it by adjusting the selection window size. Experimental results validate the efficiency of the proposed method.
Index Terms:
5G NR V2I, SPS, Fairness Access.
I Introduction
With the 3rd Generation Partnership Project (3GPP) Release 16, the initial Vehicle-to-Everything (V2X) standard grounded in the 5G New Radio (NR) was introduced as a supplement to Long-Term Evolution (LTE) V2X communication[1]. NR V2X supports two types of communication: mode-1 (centralized) and mode-2 (distributed)[2, 3]. Vehicles may operate within network coverage in mode-1, and resources are scheduled through the Base station(BS)[4, 5]. In contrast, mode-2 allows vehicles to autonomously allocate resources with the sensing-based semi-persistent scheduling (SPS) mechanism, significantly enhancing the flexibility of resource scheduling for vehicles[6]. Modern vehicles are often outfitted with cameras, LiDARs, and other sensors to sense the environment[7]. However, due to limited onboard computational resources, it is challenging for vehicles to process the large-scale data they produce[8]. To extract useful information, vehicles often rely on cloud computing by offloading complex computational tasks to a central cloud , which processes the data and returns the results[9]. However, traditional cloud computing is unsuitable for high-speed vehicular scenarios, as long-distance transmission introduces significant communication delays, which can be dangerous in mobile vehicular environments[10, 11]. To address this, Vehicle-to-Infrastructure (V2I) technology is applied in vehicular network scenarios to obtain real-time data that assists with driving[12]. In NR V2I mode-2, vehicles typically employ sensing-based SPS mechanisms for resource scheduling[13]. However, vehicles on different lanes often travel at varying speeds, resulting in different durations of time spent within an RSU’s coverage area[14]. Furthermore, transmission failures may occur, leading to an unequal amount of successfully transmitted data for vehicles of different speeds[15]. This difference, referred to as unfair access, often results in high-speed vehicles receiving less information than low-speed vehicles, which increases the likelihood of incorrect decisions and safety risks[16].
In summary, designing an access scheme that ensures fair data access in vehicular networks is of great importance. As far as we know, there has been no prior research specifically addressing data access fairness in 5G NR V2I, which motivates our work.
The key contributions of this work are summarized as below [^1] Source code can be found at https://github.com/qiongwu86/Enhanced-SPS-Velocity-adaptive-Scheme-Access-Fairness-in-5G-NR-V2I-Networks :
  • 1)
    We propose a speed-adaptive selection window adjustment scheme for sensing-based SPS scheduling in 5G NR V2I mode 2 to achieve fair network access.
  • 2)
    We define a fairness index to represent the fairness of data access for vehicles at different speeds, proving that this index is a function of both speed and the selection window size.
  • 3)
    We adopt the Non-dominated Sorting Genetic Algorithms-II (NSGA-II)[17] to solve the proposed multi-objective optimization problem. Experimental results validate the efficiency of the proposed method.
The structure of the rest of this work is outlined below: Section II provides an overview of related studies. Section III presents the system model. Section IV presents the fairness metric. Section V formulates the multi-objective optimization problem and describes its solution using the NSGA-II. Section VI details the experimental results. Section VII concludes the paper.
Figure 1: System Model
II Related Work
In this section, we will review the related works relevant to this paper.
II-A Sensing Based Semi-Persistent Scheduling
Some studies have been worked on the SPS within NR V2X Mode-2. In [18, 19], Daw et al. proposed a priority-based SPS scheme that categorizes emergency vehicles separately and introduced a complementary probabilistic collision mitigation mechanism to minimize the collision probability for high-priority vehicles in the network. In [20], Jeon et al. significantly reduced data packet conflicts in the C-V2X mode-4 by minimizing the uncertainty in resource selection for SPS. Specifically, they utilized a “lookahead” approach to eliminate collisions caused by unawareness of other users’ decisions.
II-B Fairness of network
Several studies have proposed solutions to address the fair access of vehicle caused by varying vehicle speeds. In [21], Wan et al. proposed modifying the contention window in the IEEE 802.11p protocol to address vehicle access fairness, while jointly considering the network-wide age of information to minimize latency. In [22], Harigovindan et al. employed proportional fairness (PF) in multi-rate, multi-lane V2I networks, improving the overall data transfer efficiency. In [23], Wu et al. addressed vehicle access fairness in a platooning scenario by dynamically tuning the minimum contention window according to vehicle speeds under the IEEE 802.11p.
However, these studies have not considered the fairness issue under the NR V2I Mode-2 SPS. Due to the few research focusing on vehicle access fairness in 5G NR V2X Mode 2 SPS, this motivates us to undertake this work.
III System Model
This section focuses on a detailed analysis of the system model.
III-A Scene Model
As depicted in Fig.1, we consider a highway environment featuring $N$ lanes, with an RSU deployed along the roadside as an edge server. The lanes are divided into two directions, and vehicles arrive within the RSU’s communication range following a Poisson process. Once a vehicle enters the RSU’s communication range, it will undertake information transmission with the RSU and extract useful information. For the same direction, the speed of vehicles in the same lane is the same.
Figure 2: SPS Model
III-B Sensing Based Semi-Persistent Scheduling
Each vehicle adopts 5G NR V2X mode 2 for data transmission and undergoes resource allocation through the SPS[24]. Specifically, as Fig.2, the channel is partitioned into subframes temporally, each being 1ms[25]. Within the frequency spectrum, the channel is decomposed into subchannels, each composed of multiple consecutive resource blocks (RBs). A subchannel can be divided into two parts: the side link control information (SCI)[26], occupying two RBs, and the transmission block (TB), occupying the remainder. One subchannel combined with one subframe is referred to as a physical resource block (PRB). Upon reserving a PRB, it generates a reselection counter (RC) randomly and decrements it by 1 after each transmission[27]. When $RC=0$, it will reselect resources with a probability of $1-P$; otherwise, it will continue to utilize the previous resources[28].
Vehicles autonomously select subchannels and employ the selected subchannels for transmitting TB. Simultaneously, according to the resource reservation interval (RRI)[29], they announce the subchannel chosen for transmitting the next TB to avoid other vehicles choosing the same subchannel. During the reselection process, vehicle initially identifies resources within the selection window,[30] where the selection window’s size $w$ is determined independently by the vehicle. Subsequently, the vehicle determines the resources that should be excluded, namely, according to the information provided by the perception window of size 1000ms[31]. The exclusion criteria are as follows:
  • 1)
    PRBs that have already been reserved by other vehicles will be excluded.
  • 2)
    PRBs with an average RSRP over time exceeding the threshold will be excluded.
After eliminating all the PRBs that do not satisfy the conditions, the vehicle randomly selects the remaining PRBs for transmission resources.
IV Fairness Index
This section focuses on deriving the correlation between the fairness metric for vehicles at different speeds and the parameters of speed and selection window size. Additionally, we define a fairness index to quantify the network’s fairness.
IV-A Transmission rate
Achieving system fairness implies that vehicles traveling at different speeds should transmit an equal average amount of data while within the coverage area of the edge device. Therefore, we can express this as:
${E[Bit^{i}]=C},$ (1)
where $E[\cdot]$ is the averaged operation. $Bit^{i}$ represents the amount of data transmitted by vehicle $i$ within the base station’s range. $C$ is a constant, and since transmission may fail with a certain probability, we consider the expected value. Specifically, $Bit^{i}$ can be expressed as:
${Bit^{i}={{C}_{i}}\cdot{{T}_{i}}},$ (2)
Therefore, Eq.(1) can be rewritten as:
${{C}_{i}}\cdot{{T}_{i}}\cdot{{P}_{PRR}}^{{}_{i}}=C,$ (3)
where ${C}_{i}$ represents the transmission rate of vehicle $i$, and ${T}_{i}$ denotes the time vehicle $i$ spends within the BS’s range. ${{P}_{PRR}}^{{}_{i}}$ is the probability of successful transmission. Therefore, ${T}_{i}$ can be expressed as:
${{T}_{i}}=\frac{R}{{{v}_{i}}},$ (4)
where, $R$ represents the coverage range of the BS, and ${v}_{i}$ denotes the speed of vehicle. Considering the limits of our proposed scheme, and based on Shannon’s theorem, ${C}_{i}$ can be expressed as:
${{C}_{i}}=B{{\log}_{2}}(1+\frac{{{p}_{i}}\cdot{{h}_{i}}(t)\cdot{{({{d}_{i}}(t))}^{-\partial}}}{{{\sigma}^{2}}}),$ (5)
[…]
${{d}_{i}}(t)=\left\|{{P}_{o}}^{i}-{{P}_{o}}^{B}\right\|,$ (6)
where ${P}_{o}^{i}$ represents the position of vehicle $i$, and ${P}_{o}^{B}$ represents the position of the BS. The ${P}_{o}^{i}$ can be described as:
${{P}_{o}}^{i}(t)=({{v}_{i}}t,0,0).$ (7)
According to [32], we adopt an autoregressive (AR) model to express the correlation between ${h}_{i}(t)$ and ${h}_{i}(t-1)$:
${{h}_{i}}(t)={{\rho}_{i}}{{h}_{i}}(t-1)+e(t)\sqrt{1-\rho_{i}^{2}},$ (8)
where ${\rho}_{i}$ represents the channel correlation coefficient during successive time intervals, and $e(t)$ is a complex Gaussian random error vector. Considering the mobility of the vehicle, which introduces Doppler effects. we use Jake’s fading spectrum, ${{\rho}_{i}}={{J}_{0}}(2\pi f_{d}^{i}t)$, where ${{J}_{0}}(\cdot)$ is the zeroth-order Bessel function of the first kind. $f_{d}^{i}t$ represents vehicle $i$’s Doppler frequency. The Doppler shift can be expressed as:
$f_{d}^{i}=\frac{v_{i}}{{{\Lambda}_{0}}}\cos\theta,$ (9)
where ${\Lambda}_{0}$ represents the wave length, and $\cos\theta$ is the angle between the communication direction and vehicle’s direction of motion .
IV-B Successful decoding probability
Next, we will analyze ${{P}_{PRR}}^{{}_{i}}$, which represents the probability that the data packet transmitted by vehicle $i$ is successfully decoded by the BS:
${{P}_{PRR}}^{{}_{i}}=\prod\limits_{j\neq\text{i}}{(1-{{\delta}^{j}}_{COL})\cdot}\prod\limits_{j\neq\text{i}}{(1-{{\delta}^{j}}_{HD})},$ (10)
where ${{\delta}^{j}}_{COL}$ represents the data packet collisions’ probability between vehicle $i$ and vehicle $j$. Multiple vehicles can occupy the same Physical Resource Block (PRB). When several vehicles nearly simultaneously attempt to select resources, there is a possibility that they may choose the same PRB. Based on the model in [33], ${{\delta}^{j}}_{COL}$ can be expressed as:
${{\delta}^{j}}_{COL}={{P}_{O}}{{P}_{SH|O}}\frac{{{C}_{Ca}}}{N_{Ca}^{2}},$ (11)
where ${P}_{O}$ represents the probability of overlap between the selection windows of vehicle $i$ and vehicle $j$, and ${P}_{SH|O}$ is the probability which vehicle $i$ and vehicle $j$ select resources from their shared selection window[34]. Assuming the overlap occurs, represents the average count of candidate PRBs shared between vehicle $i$ and vehicle $j$, and $N_{Ca}$ represents the average number of candidate PRBs[35]. The candidate PRBs are the PRBs that are available for transmission after the re-selection steps (2)/(3). The sensing window size is 1000ms[w219]. ${P}_{O}$ can be mathematically expressed as:
${{P}_{O}}=\frac{{{w}_{i}}+{{w}_{j}}+1}{1000\cdot{{2}^{\mu}}RRI}.$ (12)
${P}_{SH|O}$ can be mathematically expressed as:
${{P}_{SH|O}}={{(\frac{{{N}_{Sc}}{{N}_{Sh}}}{{{N}_{r}}})}^{2}},$ (13)
where ${N}_{Sc}$ represents the number of subchannels, and ${N}_{Sh}$ is the shared resources’ number within the overlapping selection window. ${N}_{Sh}$ can be expressed as:
${{N}_{Sh}}=\frac{({{w}_{i}}+1)({{w}_{j}}+1)}{{{w}_{i}}+{{w}_{j}}+1}.$ (14)
${{\delta}^{j}}_{HD}$ represents the probability that vehicles use the same time slot for transmission. Due to the half-duplex nature of vehicle communication[36], the receiver cannot decode the data packet from the transmitter if both vehicles transmit simultaneously, leading to a transmission failure. Based on the model in [33], this can be expressed as:
${{\delta}^{j}}_{HD}=\frac{{{\tau}_{j}}}{1000},$ (15)
where ${\tau}_{j}$ represents the packet generation frequency of vehicle $j$. Therefore, it is known that ${{P}_{PRR}}^{{}_{i}}$ is a function of $w$, where $w$ typically refers to the resource selection window.
IV-C Fairness index
For brevity, we only consider the fairness index at a certain time. Based on the above analysis, Eq.(3) can be further expressed as:
$C=$ $B{{\log}_{2}}(1+\frac{{{p}_{i}}\cdot{{h}_{i}}\cdot{{({{d}_{i}})}^{-\partial}}}{{{\sigma}^{2}}})\cdot\frac{R}{{{v}_{i}}}$ (16)
$\cdot\prod\limits_{j\neq\text{i}}{(1-{{\delta}^{j}}_{COL})\cdot}\prod\limits_{j\neq\text{i}}{(1-{{\delta}^{j}}_{HD})}.$
We eliminate the items that have no relation to vehicle $i$. Therefore, Eq. (16) can be further expressed as:
$K_{index}^{i}=\frac{C}{{{C}^{\prime}}}=$ ${{\log}_{2}}(1+\frac{{{p}_{i}}\cdot{{h}_{i}}\cdot{{({{d}_{i}})}^{-\partial}}}{{{\sigma}^{2}}})$ (17)
$\cdot\frac{\prod\limits_{j\neq\text{i}}{(1-{{\delta}^{j}}_{COL})}}{{{v}_{i}}},$
where,
${C}^{\prime}=B\cdot R\cdot\prod\limits_{j\neq\text{i}}{(1-{{\delta}^{j}}_{HD})}.$ (18)
Thus, we have derived the fairness index for vehicle $i$. Furthermore, since ${d}_{i}$ is a function of $v_{i}$ and ${{P}_{PRR}}^{{}_{i}}$ is a function of $w$, the fairness index $K_{index}^{i}$ is a function of both $v$ and $w$. Therefore, by knowing the vehicle’s speed, we can adaptively adjust the vehicle’s selection window based on speed to achieve overall network fairness.
By averaging the speed and window size of all vehicles in the network, we can obtain:
$K_{index}={{\log}_{2}}(1+\frac{{{p}_{i}}\cdot{{h}_{i}}\cdot{{({{d}_{i}}(\bar{v}))}^{-\partial}}}{{{\sigma}^{2}}})\cdot\frac{\prod\limits_{j\neq\text{i}}{(1-{{\delta}^{j}}_{COL}(\bar{w}))}}{{{{{v}}}_{i}}}$ (19)
where $\bar{v}$ represents the network’s average speed, and $\bar{w}$ represents the network’s average window size. The fairness index $K_{index}$ can measure the overall network’s fairness. The vehicle is achieving fair access when $K_{index}^{i}$ approaches $K_{index}$.
V Optimization Problem and Solution
In this section, we formulate a multi-objective optimization problem and employ the NSGA-II algorithm [17]. As a result, we can derive the optimal selection window size that achieves network fairness.
V-A Optimization Objective
The optimization goal is to adjust each vehicle’s selection window sizes to ensure that the data transmitted between each vehicle and the RSU is similar. This implies that the fairness index $K_{index}^{i}$ approaches the network’s average fairness index $K_{index}$. Accordingly, the optimization objective functions can be formulated as:
Objective 1 to $\bm{N}$: Minimize the variation between the fairness index on different lanes and network’s fairness index.
$\min_{\bm{w}}$ $\bm{F}(\bm{w})=[F_{K_{1}}(\bm{w}),F_{K_{2}}(\bm{w}),\dots,F_{K_{N}}(\bm{w})]^{T}$ (20)
$s.t$
$w^{LB}\leq w^{i}\leq w^{UB},i\in[1,\ldots,N],$
where
$F_{K_{i}}(\bm{w})=\left|K_{index}(\bm{w})-K_{index}^{i}(\bm{w})\right|,i\in[1,\ldots,N],$ (21)
$\bm{w}=\{w^{1},w^{2},...,w^{N}\}$. $w^{LB}$ and $w^{UB}$ indicate the minimum and maximum bounds of the selection window sizes, as defined in the NR V2I SPS based on [6].
Thus, we have defined a multi-objective optimization problem, which enables achieving overall network fairness by adjusting the selection window sizes of individual vehicles.
V-B Optimization Solution
Input: $\bm{v}=\{v_{1},v_{2},\dots,v_{N}\}$, $N_{max}$: Maximum number of generations, $M$: population size, Threshold
Output: optimal solution $\bm{w}^{*}$
1 Initialize the population $P_{0}=\{\bm{w_{1}},\bm{w_{2}},\dots,\bm{w_{M}}\}$ within $[w^{LB},w^{UB}]$ randomly;
2 for $n=0$ to $N_{max}$ do
3       Crossover: Generate offspring $P_{n}^{C}$ from $P_{n}$ using crossover;
4       Mutation: Apply mutation to $P_{n}^{C}$ to produce $P_{n}^{M}$;
5       Population Merge: Form $Q_{n}=P_{n}\cup P_{n}^{M}$;
6       Objective function value: Calculate the value of objective function $F_{K}(\bm{w_{i}})=\{f_{1}(\bm{w_{i}}),f_{2}(\bm{w_{i}}),\dots,f_{M}(\bm{w_{i}})\}$ based on Eq.(21), for all $\bm{w_{i}}\in Q_{n}$;
7       Non-dominated Sorting: Compute $\bm{F}=\{F_{1},F_{2},\dots\}$ for $Q_{n}$ with $F_{K}(\bm{w_{i}})$;
8       Crowding Distance: Compute crowding distance $d(\bm{w_{i}})$ for all $\bm{w_{i}}\in Q_{t}$ with $F_{K}(\bm{w_{i}})$;
9       Selection: Select the top $M$ individuals in $Q_{n}$ as $P_{n+1}$ using $\bm{F}$ and $d(\bm{w_{i}})$;
10
11Optimization Phase: ;
12 $P_{\text{filtered}}=\bm{w}\in$$P_{n+1}$ : $F_{K_{i}}(\bm{w})\leq\text{Threshold},\forall i$;
13 $\bm{w}^{*}=\arg\min_{\bm{w}\in P_{\text{filtered}}}\sum_{i=1}^{N}F_{K_{i}}(\bm{w})$;
14
return $\bm{w}^{*}$
Algorithm 1 Non-dominated Sorting Genetic Algorithm II
We use the NSGA-II algorithm [17] to address the optimization problem. Each population consists of $N$ individuals, representing the vehicle selection windows. The population size is $M$. The algorithm can be organized into three parts: initialization phase, iteration phase, and optimization phase. The algorithm’s procedure is presented in Algorithm 1.
1) Initialization Phase: We randomly initialize a set of selection windows within the range $[w^{LB},w^{UB}]$.
2) Iteration Phase: Firstly, crossover and mutation are performed. Crossover involves exchanging parts of two parent individuals’ genes to generate new offspring, while mutation randomly modifies an individual’s genes to prevent the algorithm from converging to a local optimum. At this point, the parent and offspring populations are merged into a new population $Q_{n}$ for the subsequent selection operation. Next, we use $\bm{v}=\{v_{1},v_{2},\dots,v_{N}\}$ to calculate the objective function values $F_{K}(\bm{w})$ according to Eq.(21). Then we perform non-dominated sorting. Specifically, if individual $A$ is equal to or better than individual $B$ in every objective, and better than $B$ no fewer than one objective, individual $A$ is regarded as dominate individual $B$. Non-dominated sorting divides the population into multiple levels according to each individual’s dominance rank. Thus we derive $\bm{F}=[F_{1},F_{2},\dots]$, $F_{1}$ means the individuals which rank first, and then, crowding distance is calculated as follow[37]:
$d(\bm{w_{i}})=\sum_{j=1}^{N}\left(\frac{f_{j}^{i+1}(\bm{w_{i}})-f_{j}^{i-1}(\bm{w_{i}})}{f_{j}^{max}-f_{j}^{min}}\right),$ (22)
where $f_{j}^{i+1}$ and $f_{j}^{i-1}$ respectively represent the objective value on objective $j$ for the previous and next individual of the dominance rank individual $i$. $f_{j}^{max}$ and $f_{j}^{min}$ respectively represents the maximum and minimum objective values of all individuals on the objective $j$[38].
Then, parent individuals are selected based on non-dominated sorting and crowding distance. The selection prefers individuals that are ranked higher in non-dominated sorting, and within the same rank, individuals with larger crowding distances are preferred. Thus, the merged population $Q_{n}$ undergoes non-dominated sorting and selection. Finally, $M$ individuals with higher domination rank and larger crowding distances are selected to form the new population $p_{n+1}$.
3) Optimization Phase: In this phase, the optimal population is selected from the Pareto front. The optimal population must ensure that the difference between the fairness index of each vehicle is smaller than a threshold, while also minimizing the total objective function values.
Thus, the optimal selection window $\bm{w}^{*}$ is obtained.
VI Numerical Simulation And Analysis
In this section, we validate the effectiveness of our proposed scheme. The simulation experiments are implemented using Python 3.9. The scenario is set as a two-way highway with four lanes, where the speed limit is between 20 and 30 m/s, and the speed difference between adjacent lanes is within 4 m/s. The window size limits and time slot length are configured according to the 5G NR V2I Mode 2 settings.
Figure 3: Performance Indicators of the NSGA-II Algorithm
Figure 4: Optimal Selection Window Versus Average Velocity
Figure 5: Optimal Selection Window Versus Average Velocity
Fig. 3 presents the performance metrics of the NSGA-II for solving the proposed multi-objective problem, evaluated using hypervolume (HV), Inverted Generational Distance (IGD), Generational Distance (GD), and Spacing.
The hypervolume (HV) measures the extent to which the solution set covers the objective space, progressively increasing and stabilizing, indicating high diversity. Generational Distance (GD) evaluates the proximity of the obtained solutions to the true Pareto front, with values steadily decreasing and stabilizing, demonstrating effective convergence. Inverted Generational Distance (IGD) considers both diversity and convergence, yielding higher values than GD. Lastly, Spacing measures the uniformity of solution distribution, remaining consistently low and indicating an even spread of solutions.
Fig. 4 illustrates the optimal window size variation for three vehicles traveling in different lanes as the average vehicle speed increases. It can be observed that vehicles with higher speeds tend to have smaller optimal window sizes. This is because higher speeds reduce the transmission time with RSU, necessitating a decrease in the window size to maximize data transmission. Moreover, as the average network speed increases, the window sizes for all vehicles decrease. This is also attributed to the challenge of achieving fairness at higher speeds, which requires a reduction in the window size to balance data transmission across vehicles effectively.
In Fig. 5, we present a comparison of the objective values when vehicles adopt the optimal selection window versus the standard selection window. we can see that when using the standard selection window, the objective values of vehicles progressively increase, indicating a gradual loss of fairness in access. This phenomenon arises because faster vehicles will have less staying time in RSU coverage, resulting in transmitting less data, leading to unfairness. However, when the optimal selection window is employed, both the rate of increase and the initial values are significantly smaller compared to the standard window. This improvement is attributed to the vehicles’ ability to adaptively adjust the optimal window based on their speed, thereby striving to maintain fairness.
VII Conclusion
In this paper, we considered the fairness in data access within vehicular networks and proposed a multi-objective optimization method based on adjusting the selection window of the SPS in NR V2I. The goal is to ensure fair access to the base station for vehicles at varying speeds. The simulation results lead to the following conclusions:
  • Vehicle speed significantly impacts access fairness. Under the same conditions, higher vehicle speeds result in reduced data exchange with the RSU, making it more challenging to achieve fairness.
  • The optimal selection window decreases progressively with increasing vehicle speed to reduce data transmission time costs, thereby improving fairness.
In future research, optimization can be broadened to simultaneously consider the minimization of age of information.